# Database Views as Explanations for Relational Deep Learning

### v1

@@ Abstract.
We investigate the realization of the framework in the case of hetero-GNNs. We develop heuristic algorithms that avoid the exhaustive search over the space of all databases. We propose techniques that are model-agnostic, and others that are tailored to hetero-GNNs via the notion of learnable masking. Our approach is evaluated through an extensive empirical study on the RelBench collection, covering a variety of domains and different record-level tasks. The results demonstrate the usefulness of the proposed explanations, as well as the efficiency of their generation.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at URL_TO_YOUR_ARTIFACTS.
@@ 1. Introduction
Machine learning over relations was traditionally done through a reduction to learning over unstructured data: flatten the data into a feature matrix, with a collection of predefined queries that produce features for every prediction instance, and then learn a flat model (Perlich and Provost, 2003; Kimelfeld and Ré, 2018; Bakshi et al., 2021). Progress in artificial neural networks, and specifically Graph Neural Networks (GNNs) (Wu et al., 2020) and graph transformers (Müller et al., 2024) changed this practice: the best performing models now deploy prediction models directly from the structured data without ad-hoc feature extraction(Chen et al., 2025; Wang et al., 2024).
Specifically, deep learning is commonly applied to relational databases by translating them into a graph representation: nodes represent database entities (usually tuples), and edges represent relationships (usually foreign-primary key constraints). GNNs or graph transformers are then trained on these graphs (Fey et al., 2024; Cappuzzo et al., 2020; Tönshoff et al., 2023; Schlichtkrull and Søgaard, 2017). To handle the varying types of nodes and edges determined by the database schema, heterogeneous graph models are used (Zhao et al., 2021; Wang et al., 2019; Chen et al., 2025; Wang et al., 2024). This creates a two-paradigm pipeline: the relational database defines the graph structure, while the (graph-based) machine learning library learns the model. At inference time, new data is exported to the graph representation, predictions are computed from the learned model, and results are reintegrated into the database.
@@
The neural network model induced by the above graph-learning approach over the database can be viewed as a complex numerical function constructed from the database’s tuples and values, and parameterized by a large number of learned weights. Consequently, as in other modalities of deep learning, the model’s internal behavior is inherently opaque and cannot be readily understood through direct inspection. To address this, the field of Explainable Artificial Intelligence (XAI) provides tools that shed light on different aspects of the model (Angelov et al., 2021; Arenas, 2024; Marques-Silva, 2023; Darwiche, 2023). These include local explanations, which account for the model’s output on a specific instance, and global explanations, which aim to characterize the model’s behavior as a whole. Explanations may be contrastive (sometimes called counterfactual (Barceló et al., 2025)), identifying hypothetical changes that would alter the model’s outcome, or abductive, pointing to focused parts of the input that suffice to determine the outcome (Marques-Silva, 2023).
In this work, we study global abductive explanations for deep relational models. Our focus is on settings where the underlying schema consists of multiple interrelated tables with numerous attributes—a common scenario in real-world databases. We aim to identify concise, informative subsets of the data that best account for the model’s predictions. Specifically, we address the question: Which parts of the database does the model rely on most when making predictions? Hence, our goal is to provide global insights into the model’s behavior by uncovering the database fragments that are most influential across predictions.
@@
To illustrate the nature of our explanations, consider the “rel-trial” database (Initiative et al., 2019) from the RelBench collection (Robinson et al., 2024). Suppose we wish to explain a binary classifier that solves the following task: “Predict whether a trial will achieve its primary outcome.” Here, a trial corresponds to a record in the relation Studies. Our approach identifies only $\sim 10\%$ out of $140$ columns together with the $10$ relevant primary-foreign key pairs, which capture the schema components that largely determine the model’s output. In particular, our explanation highlights the attributes name, city, and country from the Facilities relation, which is linked to Studies via the intermediate relation Facilities_studies. This indicates that the names and locations of the utilized facilities highly influence the success or failure of the study. Interestingly, no attributes from the Outcomes relation, e.g., the outcome type or description, are included in the explanation, indicating that they are not actually as useful for the model to predict whether this outcome is achieved. Our explanations further show that specific study designs (i.e., records in the Designs relation) that utilize simple interventions are important for the model’s prediction. A detailed case study follows in Section 6.3. To verify that the schema components identified by our explanation capture the most important parts of the database, we train a smaller model using only those components, which achieves similar performance (ROC_AUC drops only slightly from $71.28$ to $69.78$). This is still high performance for the task (see (Robinson et al., 2024) for reference), while reducing the database by $\sim 80\%$.
@@
To establish insights such as those in Example 1, we propose a general framework in which explanations are expressed as view definitions written in SQL (or any other query language of choice). This grounds explanations in a representation that is familiar to database users, in contrast to common alternatives that require machine learning expertise to interpret (Räuker et al., 2023). For graph-based models, this would require knowledge of the graph modeling details or even the inner workings of GNNs (Yuan et al., 2022).
@@
To this aim, we deploy the classical notion of determinacy, introduced by Nash, Segoufin, and Vianu (Nash et al., 2010), that captures in a fundamental way when a collection of database views is enough to determine the result of a query. For reasons of usability, flexibility, and practical implementation, we propose a soft and statistical adaptation of this concept. Roughly speaking, the definition we propose identifies an explanation as accurate if the model is likely to have similar predictions on our database as it would have on alternative databases, as long as they agree on the views. Figure 1 shows a visual illustration of the concept.
In addition to accuracy, we seek concise explanations that are easy to digest. Clearly, the verbosity of an explanation depends on both the number of views and the complexity of each view. Our framework naturally supports the tradeoff between accuracy and conciseness, allowing users to balance interpretability with fidelity to the model’s behavior.
Nevertheless, the theoretical definition that we propose for explanations does not lend itself directly to an implementation. We develop heuristic techniques for the case of GNNs and important classes of views: projections, joins over foreign keys, and selections (filters). These heuristics are generally based on learning masked variations of the GNN at hand, where we place the masks in components of the GNN that correspond to the different kinds of views.
@@
Lastly, we present an empirical evaluation of our techniques on the RelBench collection of databases and node-level tasks (Robinson et al., 2024). In particular, the experiments show the tradeoff between the accuracy and conciseness of the explanations. As a side benefit, they also show that, for several of the RelBench tasks, we can achieve models of similar quality to the ones publicly available, while using only a fraction of the database.
@@
Our main contributions. 1 We introduce a model-agnostic framework for generating relational explanations of learned model behavior over relational databases, based on a statistical adaptation of determinacy. The framework uses relational queries, supports user-controlled granularity, and does not rely on internal model representations. 2 We instantiate the framework for relational deep learning using GNNs, and develop an efficient masking-based heuristic to generate explanations without explicitly estimating soft determinacy. 3 We conduct extensive experiments on the RelBench benchmark across several databases and tasks, proposing evaluation baselines and metrics. Our method consistently produces high-quality explanations at significantly lower computational cost.
@@ 2.1. Relational Database Concepts
Databases
A relation schema $\mathcal{R}$ is associated with a sequence $\mathsf{att}(\mathcal{R})=(A_{1},\dots,A_{k})$ of distinct attributes, and a key signature $\mathsf{key}(\mathcal{R})=(B_{1},\dots,B_{\ell})$, so that $\mathsf{key}(\mathcal{R})$ is a subsequence of $\mathsf{att}(\mathcal{R})$. Each attribute $A_{i}$ is associated with a (finite or infinite) domain, denoted $\mathsf{dom}(A_{i})$. If $\mathord{\mathbf{C}}=(C_{1},\dots,C_{m})$ is a sequence of attributes of $R$, then we denote by $\mathsf{dom}(\mathord{\mathbf{C}})$ the set $\mathsf{dom}(C_{1})\times\dots\times\mathsf{dom}(C_{m})$. A tuple $t$ over a relation schema $\mathcal{R}$ is a member of $\mathsf{dom}(\mathsf{att}(R))$. For $i=1,\dots,k$, we denote by $t[A_{i}]$ the $i$-th value in $t$ (corresponding to the attribute $A_{i}$). If $\mathord{\mathbf{C}}=(C_{1},\dots,C_{m})$ is a sequence of attributes from $\mathsf{att}(\mathcal{R})$, then we denote by $t[\mathord{\mathbf{C}}]$ the tuple $(t[C_{1}],\dots,t[C_{m}])$.
@@
Relational queries and predictive models
Let $\mathcal{S}$ be a database schema. A query $Q$ over $\mathcal{S}$ is associated with a relation schema, which we consistently denote by $\mathcal{R}_{Q}$, and it maps every database $D$ over $\mathcal{S}$ to a relation $Q(D)$ over $\mathcal{R}_{Q}$.
@@
  • existence of prediction attribute $A$: $\mathsf{att}(\mathcal{R}_{M})=(\mathsf{key}(R),A)$, where $A$ is an attribute with $\mathsf{dom}(A)=\mathbb{R}^{d}$, where $\mathbb{R}$ stand for the set of real numbers.
@@
Let $M$ be an $R$-model and $D$ a database, both over the schema $\mathcal{S}$. Let $s$ be an instance (i.e., a key tuple of $D(R)$) to which we apply the model We denote by $M_{D}(s)$ the unique prediction of $M$ for $s$, that is, the value $t[A]\in\mathbb{R}^{d}$ of the unique tuple $t\in M(D)$ with $t[\mathsf{key}(R)]=s$.
Determinacy
We briefly recall the concept of determinacy by Nash et al. (2010). Let $\mathcal{S}$ be a schema, let $V_{1}$, …, $V_{m}$ be queries over $\mathcal{S}$, referred to as views, and let $Q$ be a query over $\mathcal{S}$. We say that $\mathord{\{V_{1},\dots,V_{m}\}}$ determines $Q$ if for all databases $D_{1}$ and $D_{2}$, if $V_{i}(D_{1})=V_{i}(D_{2})$ for every $i=1,\dots,m$, then $Q(D_{1})=Q(D_{2})$. Hence, knowing the result of every view suffices to determine the result of $Q$ without even looking at the database itself.
@@ 2.2. Heterogeneous Graphs
Hetero-graphs
A heterogeneous featured graph, or just hetero-graph for short, is a quadruple $\mathord{\mathcal{G}}=(V,E,\tau,\mathord{\mathbf{X}})$ where:
@@
Node-centric models
Let $\mathord{\mathcal{G}}=(V,E,\tau,\mathord{\mathbf{X}})$ be a hetero-graph, and let $\alpha\in\tau(V)$ be a type, which we refer to as a prediction entity type. A $\alpha$-model (of dimension $d$) is a function $M:V[\alpha]\rightarrow\mathbb{R}^{d}$, mapping every $\alpha$-node $v$ to a predicted vector $M(v)$. Note that the dimension $d$ depends on the task (e.g., it is one in binary classification and regression, and $k$ in $k$-class classification), and is not necessarily the same as $d_{\alpha}$.
@@ 2.3. Heterogeneous Graph Neural Networks
A GNN defines a model over a graph by applying message passing to produce a vector representation of each node; in each step, called layer, the vector of a node is updated by aggregating the vectors of its neighbors, referred to as messages. A hetero-GNN operates over a hetero-graph similarly, except that messages of each edge type are aggregated separately. We give a more precise definition of a typical hetero-GNN next. We note that our definition is slightly more general and detailed than common definitions (e.g., (Barceló et al., 2022)), stressing the role of labels and directionality in the GNN.
@@
for every node $v$, edge label $\lambda$ and orientation $o\in\mathord{\{{\mathsf{in}},{\mathsf{out}}\}}$ such that $\mathcal{N}^{o}_{\lambda}(v)$ is defined.
Here, $\mathsf{AGG}$ is an aggregation function such as average or max, and $\mathord{\{\mskip-5.0mu\{\cdot\}\mskip-5.0mu\}}$ is the bag (multiset) notation. Next, for $\ell>0$ we define
@@
for some ordering $(\lambda_{1},o_{1}),\dots,(\lambda_{q},o_{q})$ of all pairs $(\lambda,o)$ where $\mathcal{N}^{o}_{\lambda}(v)$ is defined. Here, $\mathsf{UPD}$ is an update function, which is realized as a neural model. A typical example of model architecture for hetero-GNNs (Schlichtkrull et al., 2018) involves a linear combination followed by an activation function (e.g., ReLU). The model $M$ is given by
$M(v)=\mathsf{MLP}(\mathord{\mathbf{h}}^{(L)}(v))$
where $\mathsf{MLP}$ is a multilayer-perception function that transforms the last layer into the actual prediction. The learnable parameters are incorporated in the functions $\mathsf{AGG}$, $\mathsf{UPD}$, and $\mathrm{MLP}$.

### v2

@@ Abstract.
We investigate the realization of the framework in the case of hetero-GNNs, and develop a model-specific approach via the notion of learnable masks. For comparison, we propose model-agnostic heuristic baselines and show that our approach is both more efficient and achieves better explanation quality in most cases. Our extensive empirical evaluation on the RelBench collection across diverse domains and record-level tasks demonstrates both the usefulness of our explanations and the efficiency of their generation.
@@ 1. Introduction
Machine learning over relations was traditionally done through a reduction to learning over unstructured data: flatten the data into a feature matrix, with a collection of predefined queries that produce features for every prediction instance, and then learn a flat model (Perlich and Provost, 2003; Kimelfeld and Ré, 2018; Bakshi et al., 2021). Progress in deep learning, and specifically Graph Neural Networks (GNNs) (Wu et al., 2020) and graph transformers (Müller et al., 2024) changed this practice: the best performing models now deploy prediction models directly from the structured data without ad-hoc feature extraction (Chen et al., 2025; Wang et al., 2024).
Specifically, deep learning is commonly applied to relational databases by translating them into a graph representation: nodes represent database entities (usually tuples), and edges represent relationships (usually foreign-primary key constraints). GNNs or graph transformers are then trained on these graphs (Fey et al., 2024; Cappuzzo et al., 2020; Tönshoff et al., 2023; Schlichtkrull and Søgaard, 2017). To handle the varying types of nodes and edges determined by the database schema, heterogeneous graph models are used (Zhao et al., 2021; Wang et al., 2019; Chen et al., 2025; Wang et al., 2024). This creates a two-paradigm pipeline: the relational database defines the graph, while the (graph-based) machine learning library learns the model. At inference time, new data is exported to the graph representation, predictions are computed from the learned model, and results are reintegrated into the database.
@@
The deep model induced by the above graph ML approach over the database can be viewed as a complex numerical function constructed from the database’s tuples and values, and parameterized by learned weights. Consequently, the model’s behavior is inherently opaque and cannot be readily understood through direct inspection. To address this, the field of Explainable Artificial Intelligence (XAI) provides tools that shed light on different aspects of the model (Angelov et al., 2021; Arenas, 2024; Marques-Silva, 2023; Darwiche, 2023). These include local explanations, which account for the model’s prediction on a specific instance, and global explanations, which aim to characterize the model’s general behavior. Explanations may be contrastive (or counterfactual (Barceló et al., 2025)), identifying hypothetical changes that would alter the model’s outcome, or abductive, pointing to focused parts of the input that suffice to determine the outcome (Marques-Silva, 2023).
In this work, we study global abductive explanations for deep relational models. Our focus is on settings where the underlying schema consists of multiple interrelated tables with numerous attributes—a common scenario in real-world databases. We aim to identify concise subsets of the data that best account for the model’s predictions. Specifically, we address the question: Which parts of the database does the model rely on most when making predictions? Hence, our goal is to provide global insights into the model’s behavior by uncovering the database fragments that are most influential across multiple predictions.
@@
To illustrate the nature of our explanations, consider a relational database with schema R(rid, $a_{1}$, $\ldots$, $a_{5}$), S(sid, rid, $b_{1}$, $\ldots$, $b_{5}$, flag), T(tid, rid, $c_{1}$, $\ldots$, $c_{5}$), where rid is the primary key of $R$ and a foreign key in both $S$ and $T$. All attributes $a_{i},b_{i},c_{i}$ are numerical, while flag is boolean. Consider two binary classification tasks predicting labels for $R$-records. The ground truth of Task $1$ corresponds to whether at least $3$ linked $S$-records have flag = TRUE, while Task $2$ requires only at least $3$ linked $S$-records (regardless of flag values). Task $2$ thus depends purely on the join structure (on R.rid = S.rid), while Task $1$ additionally depends on a specific attribute (i.e., the flag).
Now suppose a data scientist has trained a relational deep learning model that achieves high accuracy (ROC-AUC ¿ 0.95) on both tasks. Applying our approach, the data scientist can understand what drives each model’s predictions. For Task 1, explanations focused on important features reveal that only S.flag is critical, while all other feature columns ($a_{i},b_{i},c_{i}$) are excluded. To verify whether the relational structure also contributes, the data scientist requests relationship-focused explanations, which show that only the $R$-$S$ join is significant (while the $R$-$T$ join plays no role). Together, these findings show the model aggregates flag values over $S$-records linked to each $R$-record. Switching to Task 2, feature-focused explanations identify no feature column as important. How can the model predict accurately without features? Our data scientist now pivots to relationship-focused explanations and finds that the $R$-$S$ join is important. This reveals that the model relies purely on connectivity structure (i.e., counting linked records) rather than attribute values.
This workflow, where users can select explanation types tailored to databases (columns, relationships, etc.) distinguishes our framework from conventional XAI approaches (Breiman, 2001; Ribeiro et al., 2016; Lundberg and Lee, 2017; Sundararajan et al., 2017). In our example, traditional feature importance methods would fail to convey that Task 2 depends solely on connectivity structure. We implemented these tasks and confirmed our method produces these explanations.
@@
To gain insights such as those in Example 1, we propose a framework in which explanations are expressed as view definitions in SQL (or any query language of choice). This grounds explanations in a representation familiar to database users, in contrast to common alternatives that require machine learning expertise to interpret (Casper et al., 2023). For graph ML, this would require knowledge of the graph modeling details or even the inner workings of GNNs (Yuan et al., 2023).
@@
To this aim, we deploy the classical notion of determinacy, introduced by Nash, Segoufin, and Vianu (Nash et al., 2010), that captures in a fundamental way when a collection of database views is enough to determine the result of a query. For reasons of usability, flexibility, and practical implementation, we propose a soft and statistical adaptation of this concept. Roughly speaking, the definition we propose identifies an explanation as accurate if the model is likely to have similar predictions on our database as it would have on alternative databases, as long as they agree on the views. Figure 1 shows a visual illustration. In addition to accuracy, we seek concise explanations that are easy to digest. Our framework naturally supports the tradeoff between accuracy and conciseness, allowing users to balance interpretability with fidelity to the model’s behavior.
While our framework is model-agnostic, as it applies to any relational predictive model, efficient computation requires heuristics to avoid exhaustive search over the space of candidate explanation views, with model-specific algorithms offering further advantages by leveraging model structure. We instantiate the framework for the case of hetero-GNNs, for which we propose a model-specific approach for important classes of views: projections, joins over foreign keys, and selections (filters). The approach is based on learning masked variations of the GNN at hand, by applying masks on GNN components that correspond to the different view classes.
@@
Lastly, we present an empirical evaluation of our techniques on the RelBench collection of databases and tuple-level tasks (Robinson et al., 2024). In particular, the experiments show the tradeoff between explanation accuracy and conciseness. As a side benefit, we show that, for most RelBench tasks, we can achieve models of similar quality to the ones publicly available, while using only a fraction of the database.
@@
Our main contributions. 1 We introduce a model-agnostic framework for generating relational explanations of predictive models over relational databases, based on a statistical adaptation of determinacy. The framework uses relational queries, supports user-controlled granularity, and does not rely on internal model representations. 2 We instantiate the framework for hetero-GNNs, and develop an efficient model-specific approach via learnable masking that generates explanations without explicitly estimating soft determinacy. 3 We conduct extensive experiments on the RelBench benchmark across several databases and tasks. To facilitate comparison, we implement model-agnostic heuristic alternatives and demonstrated that our model-specific method consistently produces high-quality explanations at significantly lower computational cost.
@@ 2.1. Relational Database Concepts
Databases. A relation schema $\mathcal{R}$ is associated with a sequence $\mathsf{att}(\mathcal{R})=(A_{1},\dots,A_{k})$ of distinct attributes, and a key signature $\mathsf{key}(\mathcal{R})=(B_{1},\dots,B_{\ell})$, so that $\mathsf{key}(\mathcal{R})$ is a subsequence of $\mathsf{att}(\mathcal{R})$. Each attribute $A_{i}$ is associated with a (finite or infinite) domain, denoted $\mathsf{dom}(A_{i})$. If $\mathord{\mathbf{C}}=(C_{1},\dots,C_{m})$ is a sequence of attributes of $R$, then we denote by $\mathsf{dom}(\mathord{\mathbf{C}})$ the set $\mathsf{dom}(C_{1})\times\dots\times\mathsf{dom}(C_{m})$. A tuple $t$ over a relation schema $\mathcal{R}$ is a member of $\mathsf{dom}(\mathsf{att}(R))$. For $i=1,\dots,k$, we denote by $t[A_{i}]$ the $i$-th value in $t$ (corresponding to the attribute $A_{i}$). If $\mathord{\mathbf{C}}=(C_{1},\dots,C_{m})$ is a sequence of attributes from $\mathsf{att}(\mathcal{R})$, then we denote by $t[\mathord{\mathbf{C}}]$ the tuple $(t[C_{1}],\dots,t[C_{m}])$.
@@
Relational queries and predictive models. Let $\mathcal{S}$ be a database schema. A query $Q$ over $\mathcal{S}$ is associated with a relation schema, which we consistently denote by $\mathcal{R}_{Q}$, and it maps every database $D$ over $\mathcal{S}$ to a relation $Q(D)$ over $\mathcal{R}_{Q}$.
@@
  • existence of prediction attribute $A$: $\mathsf{att}(\mathcal{R}_{M})=(\mathsf{key}(R),A)$, where $A$ is an attribute with $\mathsf{dom}(A)=\mathbb{R}^{d}$, where $\mathbb{R}$ stands for the set of real numbers.
@@
Let $M$ be an $R$-model and $D$ a database, both over the schema $\mathcal{S}$. Let $s$ be an instance (i.e., a key tuple of $D(R)$) to which we apply the model. We denote by $M_{D}(s)$ the unique prediction of $M$ for $s$, that is, the value $t[A]\in\mathbb{R}^{d}$ of the unique tuple $t\in M(D)$ with $t[\mathsf{key}(R)]=s$.
Determinacy. We briefly recall the concept of determinacy by Nash et al. (2010). Let $\mathcal{S}$ be a schema, let $V_{1}$, …, $V_{m}$ be queries over $\mathcal{S}$, referred to as views, and let $Q$ be a query over $\mathcal{S}$. We say that $\mathord{\{V_{1},\dots,V_{m}\}}$ determines $Q$ if for all databases $D_{1}$ and $D_{2}$, if $V_{i}(D_{1})=V_{i}(D_{2})$ for every $i=1,\dots,m$, then $Q(D_{1})=Q(D_{2})$. Hence, knowing the result of every view suffices to determine the result of $Q$ without even looking at the database itself.
@@ 2.2. Heterogeneous Graphs
Hetero-graphs. A heterogeneous featured graph, or hetero-graph for short, is a quadruple $\mathord{\mathcal{G}}=(V,E,\tau,\mathord{\mathbf{X}})$ where:
@@
Node-centric models. Let $\mathord{\mathcal{G}}=(V,E,\tau,\mathord{\mathbf{X}})$ be a hetero-graph, and let $\alpha\in\tau(V)$ be a type, which we refer to as a prediction entity type. A $\alpha$-model (of dimension $d$) is a function $M:V[\alpha]\rightarrow\mathbb{R}^{d}$, mapping every $\alpha$-node $v$ to a predicted vector $M(v)$. Note that the dimension $d$ depends on the task (e.g., it is one in binary classification and regression, and k1k-1 in $k$-class classification), and is not necessarily the same as $d_{\alpha}$.
@@ 2.3. Heterogeneous Graph Neural Networks
A GNN defines a model over a graph by applying message passing to produce a vector representation of each node; in each step, called layer, the vector of a node is updated by aggregating the vectors of its neighbors, referred to as messages. A hetero-GNN operates over a hetero-graph similarly, except that messages of each edge type are aggregated separately. We give a more precise definition of a typical hetero-GNN next. We note that our definition is slightly more general and detailed than common definitions (e.g., (Barcelo et al., 2022)), stressing the role of labels and directionality in the GNN.
@@
for every node $v$, edge label $\lambda$ and orientation $o\in\mathord{\{{\mathsf{in}},{\mathsf{out}}\}}$ such that $\mathcal{N}^{o}_{\lambda}(v)$ is defined. Here, $\mathsf{AGG}$ is an aggregation function such as average or max, and $\mathord{\{\mskip-5.0mu\{\cdot\}\mskip-5.0mu\}}$ is the bag (multiset) notation. Next, for $\ell>0$ we define
@@
for some ordering $(\lambda_{1},o_{1}),\dots,(\lambda_{q},o_{q})$ of all pairs $(\lambda,o)$ where $\mathcal{N}^{o}_{\lambda}(v)$ is defined. Here, $\mathsf{UPD}$ is an update function, which is realized as a neural model. A typical example of model architecture for hetero-GNNs (Schlichtkrull et al., 2018) involves a linear combination followed by an activation function (e.g., ReLU). The model $M$ is given by $M(v)=\mathsf{MLP}(\mathord{\mathbf{h}}^{(L)}(v))$ where $\mathsf{MLP}$ is a multilayer-perception function that transforms the last layer into the actual prediction. The learnable parameters are incorporated in the functions $\mathsf{AGG}$, $\mathsf{UPD}$, and $\mathrm{MLP}$.